# Boosting Visual-Language Models by Exploiting Hard Pairs

### v2

@@ 4.4 Performance of HELIP on noisy dataset
4.5 Comparison with other hard data selection method
@@
We evaluate the efficacy of the proposed method in enhancing the discriminative capacity of learned representations by comparing its zero-shot classification performance with that of other hard data mining strategies. As described in the Section 2, a common way to define hard data is through intra-modality similarity. Hence, we introduce the hard data mining methods depending on (sample level) image similarity and text similarity and denote them as IM and TM respectively. For a given target pair, we compute the cosine similarity between its image/text representation and that of the remaining dataset. The image and text representations are encoded using a pretrained Resnet50 and BERT, respectively. As a global preprocessing step, both IM and TM methods mine hard negatives. Subsequently, we integrate the mined hard negative pairs into the training pipeline of the CLIP+IM and CLIP+TM methods and optimize the original contrastive loss to finetune the model. Additionally, we employ the hard negative contrastive loss, HN-NCE, proposed by  Radenovic et al. (2023), as a baseline. HN-NCE upsamples the weight of hard-negatives identified by the current model. As shown in Table 6, when the CC3M pretrained CLIP model is combined with Helip, the performance of our pair-level hard data mining method significantly outperforms other sample-level techniques. In Figure6, we show the mined hard data that we got from three different preprocessing methods: hard pair mining (HPM), image similarity mining (IM), and text similarity mining (TM). The first row shows the image-text pairs chosen by HPM, while the second and third rows display the pairs chosen by IM and TM, respectively. We observe that the captions of the hard pairs mined with image similarity are only loosely connected with the image of the target pair. For samples mined by TM, their images are even mismatched with the caption of the target pair. The fact that pairs mined by TM are easier than IM is also reflected in Table 6, where the zero-shot performance of the CLIP+IM method consistently outperforms the CLIP+TM method across three datasets.
@@
Figure 6: HPM and fastHPM. We show the hard pairs mined by HPM and fastHPM. The quality of hard pairs mined by fastHPM is competitive with the pairs mined by HPM.
4.6 Impact of hard negative margin loss
We investigate the impact of using hard negative margin loss (HNML) on the performance of the SLIP model. In particular, our attention is directed towards an analysis of the SLIP model’s performance, which has been previously pre-trained on the CC3M dataset, when it is both further trained with HPM+HNML and left without HNML. Our approach involves a comparative analysis of the model’s zero-shot classification performance across multiple datasets including ImageNet, CIFAR 100, CIFAR 10, Caltech 101, Food 101, and Sun397. The results of our evaluation are comprehensively detailed in Table 7. These demonstrate that the SLIP model supplemented with HPM and HNML exhibits superior performance, with a performance boost of 4.51 and 3.27 compared to the SLIP and SLIP + HPM models respectively. Interestingly, the model achieved superior performance on the CIFAR 10 dataset without HNML. We postulate that this may be attributed to HNML’s ability to enhance the discriminative power of the learnt representations by employing the class distance as a cost metric. In light of this, our findings suggest that for classification datasets consisting of a larger number of subclasses, employing HNML during the training phase can lead to an increase in classification performance.
ImageNet CIFAR10 CIFAR100 Caltech101 Food101 Sun397 Avg.
SLIP 23.00 65.61 34.69 54.01 16.03 29.20 37.09
wo HNML 24.94 69.44 36.35 64.07 16.51 30.91 40.37
w HNML 26.05 68.18 37.77 66.89 17.05 33.68 41.60
Table 7: SLIP finetuned with and without hard negative margin loss. When finetuned with hard pairs, the zero-shot performance of CC3M pretrained SLIP can be further enhanced using HMNL.
4.7 Delving into hard pair mining
Impact of different encoders in HPM. We explored the effect of different pretrained encoders on HPM’s performance by alternating image and text encoders. Initially, the unsupervised pretrained DINO VITs8 (Caron et al., 2021) was paired with the SentenceT (Reimers & Gurevych, 2019) transformer, trained on over a billion internet-based sentences. This combination was later swapped for the SWAG VITb16 (Singh et al., 2022) and the T5 (Raffel et al., 2020). Additionally, experiments using OpenAI’s CLIP model (Radford et al., 2021) multimodal encoders were conducted. Interestingly, as Table 8 suggests, the encoder choice seemingly has negligible impact on HPM’s performance, likely due to the proficiency of current pretrained models in modeling intra-modal similarities. Moreover, the ability to use single-modal pretrained models and still achieve competitive or superior performance implies that there’s no assumption of having access to a high-quality CLIP model, such as OpenAI’s CLIP-400M.
@@ 4.7 Delving into hard pair mining
Performance Comparison between HPM and FastHPM. A comparison was made between the zero-shot performances of SLIP models, further trained with hard pairs obtained from both HPM and fastHPM. This comparison, summarized in Table 9, was conducted under three different settings, each maintaining the hyperparameter $k=500$. Additionally, we established subsets $\widetilde{\mathcal{D}}_{i}$ of sizes 3M and 6M, and accordingly denoted Helip with these subset sizes as Helip-3M and Helip-6M. Table 9 shows that the zero-shot performances of Helip-3M and Helip-6M remain competitive with the global HPM hard pair mining approach. These findings suggest that fastHPM offers an efficient strategy for hard pair mining, without compromising performance. Additionally, they hint at fastHPM’s potential to scale up hard pair mining in larger pre-training datasets, a promising direction for future exploration.
Imagenet CIFAR10 CIFAR100
SLIP 41.17 81.30 53.68
Helip- 3M 45.07 82.42 55.22
Helip- 6M 44.98 81.64 56.62
Helip- Full 45.64 82.31 53.79
Table 9: Zero-shot performance for SLIP + HELIP on CC12M with hard samples mined with HPM and fastHPM. Compared with hard samples mined with HPM, the fast versions are competitive with the full version.
Visual insights into HPM and FastHPM. We took the initiative to visualize the hard pairs as identified by the aforementioned three methods. Within Figure 6, the leftmost image-text pairing is earmarked as the target. The pairs in the primary row are those selected via HPM. The subsequent rows, specifically the second and third, present image-text pairings identified by the 6M fastHPM and the 3M fastHPM methods, respectively. Through a comparative visualization, it’s evident that the hard pairs pinpointed by fastHPM bear a significant resemblance to the target pair. For readers keen on delving deeper, we’ve provided an extended set of visualization outcomes in Appendix C.
Computational time analysis. Table 10 provides a comparison of the computational time required by HPM and fastHPM. The hard negative pairs preparation times listed were measured on 8 V100 GPUs, with the exception of the *** symbol, which was measured on a single V100 GPU. Given its efficiency and the performance similarities observed in Table 9, fastHPM emerges as a compelling alternative to the full HPM method.
CC3M CC12M YFCC15M
Helip- 3M - 2h18min 3h27min
Helip- 6M - 5h3min 6h19min
Helip- Full $*{}^{*}$ 9h11min 17h41min
5 Conclusion
In this study, we delve into boosting pre-trained CLIP models’ performance by more adeptly utilizing their original training dataset. This initiative arose from the recognition of the loosely connected nature of web-crawled image-text pairs, which resulted in suboptimal data utilization due to conventional CLIP loss. Our framework, Helip, introduces a cost-effective and easily integrable solution for improving existing model performance without extensive retraining or additional datasets. It selects hard pair data from their original training datasets and refines the existing models in a few epochs to immediately boost their performance. Specifically, Helip treats each text-image pair as a single point in the joint vision-language space, defining those that are close together as hard pairs. The Hard Pair Mining (HPM) strategy effectively identifies challenging hard pairs. The Hard Negative Margin Loss (HNML) was developed to improve existing models by utilizing that hard data. Empirical evaluations across various benchmarks, such as zero-shot classification, image-text retrieval, and linear probing, demonstrate the effectiveness and efficiency of Helip. We leave the discussion of future work in the appendix. For the discussion of future work, we leave it in the appendix.
Appendix A Appendix: Algorithm
Input: Hard pairs number per sample $k$ Pretrained unimodal vision model: $f_{text}$ Pretrained unimodal vision model: $f_{image}$ Dataset $\mathcal{D}=\{(x_{1}^{I},x_{1}^{T}),(x_{2}^{I},x_{2}^{T}),\cdots,(x_{N}^{I},x_{N}^{T})\}$ Threshold for visual and textual modality $\tau_{I}$ and $\tau_{T}$
for $i\in[1,N]$ do
       $\mathbf{s}\leftarrow[0,0,\cdots,0]^{\top}\in\mathbb{R}^{N}$ $I_{i}\leftarrow f_{image}(x_{i}^{I})$ $T_{i}\leftarrow f_{text}(x_{i}^{T})$ for $j\in[1,N]$ do
             $I_{j}\leftarrow f_{image}(x_{j}^{I})$ $T_{j}\leftarrow f_{text}(x_{j}^{T})$ $\vec{S}^{I}_{j}\leftarrow\frac{I_{i}\cdot I_{j}}{\left\|I_{i}\right\|_{2}\cdot\left\|I_{j}\right\|_{2}}\,\textbf{if}\,\frac{I_{i}\cdot I_{j}}{\left\|I_{i}\right\|_{2}\cdot\left\|I_{j}\right\|_{2}}>\tau_{I}\,\textbf{else}\,0$ $\vec{S}^{T}_{j}\leftarrow\frac{T_{i}\cdot T_{j}}{\left\|I_{i}\right\|_{2}\cdot\left\|T_{j}\right\|_{2}}\,\textbf{if}\,\frac{T_{i}\cdot T_{j}}{\left\|T_{i}\right\|_{2}\cdot\left\|T_{j}\right\|_{2}}>\tau_{T}\,\textbf{else}\,0$ $\mathbf{s}_{j}\leftarrow\vec{S}^{I}_{j}\cdot\vec{S}^{T}_{j}$
       end for
      $\mathcal{H}_{i}\leftarrow Argmax(\mathbf{s},k)$ if $\exists j\in\mathcal{H}_{i}$, $\mathbf{s}_{j}=0$ then
             $\mathcal{H}_{i}=\emptyset$  #  Indicate noise sample
end for
Algorithm 1 Hard Pair Mining (HPM)
Input: Hard pairs number per sample $k$ Pretrained unimodal vision model: $f_{text}$ Pretrained unimodal vision model: $f_{image}$ Dataset $\mathcal{D}=\{(x_{1}^{I},x_{1}^{T}),(x_{2}^{I},x_{2}^{T}),\cdots,(x_{N}^{I},x_{N}^{T})\}$ Threshold for visual and textual modality $\tau_{I}$ and $\tau_{T}$ Candidate pool size $C$
for $i\in[1,N]$ do
       Uniformly $C$ samples from Dataset $\mathcal{D}$, $\overline{\mathcal{D}}_{i}=\{(x_{1}^{I},x_{1}^{T}),(x_{2}^{I},x_{2}^{T}),\cdots,(x_{C}^{I},x_{C}^{T})\}$ $\mathbf{s}\leftarrow[0,0,\cdots,0]^{\top}\in\mathbb{R}^{N}$ $I_{i}\leftarrow f_{image}(x_{i}^{I})$ $T_{i}\leftarrow f_{text}(x_{i}^{T})$ for $j\in[1,C]$ do
             $I_{j}\leftarrow f_{image}(x_{j}^{I})$ $T_{j}\leftarrow f_{text}(x_{j}^{T})$ $\vec{S}^{I}_{j}\leftarrow\frac{I_{i}\cdot I_{j}}{\left\|I_{i}\right\|_{2}\cdot\left\|I_{j}\right\|_{2}}\,\textbf{if}\,\frac{I_{i}\cdot I_{j}}{\left\|I_{i}\right\|_{2}\cdot\left\|I_{j}\right\|_{2}}>\tau_{I}\,\textbf{else}\,0$ $\vec{S}^{T}_{j}\leftarrow\frac{T_{i}\cdot T_{j}}{\left\|I_{i}\right\|_{2}\cdot\left\|T_{j}\right\|_{2}}\,\textbf{if}\,\frac{T_{i}\cdot T_{j}}{\left\|T_{i}\right\|_{2}\cdot\left\|T_{j}\right\|_{2}}>\tau_{T}\,\textbf{else}\,0$ $\mathbf{s}_{j}\leftarrow\vec{S}^{I}_{j}\cdot\vec{S}^{T}_{j}$
       end for
      $\mathcal{H}_{i}\leftarrow Argmax(\mathbf{s},k)$ if $\exists j\in\mathcal{H}_{i}$, $\mathbf{s}_{j}=0$ then
             $\mathcal{H}_{i}=\emptyset$  #  Indicate noise sample
end for
Algorithm 2 fast Hard Pair Mining (fastHPM)
Input: $\mathcal{D}=\{(x_{1}^{I},x_{1}^{T}),(x_{1}^{I},x_{1}^{T}),\cdots,(x_{N}^{I},x_{N}^{T})\}$ Hard Pair Mining algorithm, $\text{HPM}()$  # or the $\text{fastHPM}()$ Pretrained unimodal vision model: $f_{text}$ Pretrained unimodal vision model: $f_{image}$ Pretrained contrastive language-image model $\{\phi_{image},\phi_{text}\}$ hyperparameters:  Hard pairs number $k$  Hard negative margin strength $\gamma$  Sampled hard negatives number $p$  Learning ratio $\eta$  Batch size $b$  Training iteration number $E$  Visual and textual modality threshold $\tau_{I}$ and $\tau_{T}$
Output: CLIP model $\{\phi_{image},\phi_{text}\}$
$\mathcal{H}\leftarrow\text{HPM}(\mathcal{D},f_{text},f_{image},k,\tau_{I},\tau_{T})$ for $iter\in[1,E]$ do
       $B\leftarrow\{z_{1},\ldots,z_{b}\}\stackrel{{\scriptstyle\text{ i.i.d. }}}{{\sim}}Uniform(\mathcal{D})$ for $z_{i}\in B$ do
             $\mathcal{H}_{i}^{p}\leftarrow\{z_{i},\ldots,z_{p}\}\stackrel{{\scriptstyle\text{ i.i.d. }}}{{\sim}}Uniform(\mathcal{H}_{i})$ $\overline{B}\leftarrow B\cup\mathcal{H}_{i}^{p}$
       end for
      Compute $\text{Compute loss}~{}\ell_{finetune}\text{, Equation (6), with samples}~{}\overline{B}$ $\phi_{image}\leftarrow\phi_{image}+\eta\cdot\partial_{\phi_{image}}\ell_{finetune}$ $\phi_{text}\leftarrow\phi_{text}+\eta\cdot\partial_{\phi_{text}}\ell_{finetune}$
end for
Appendix B Appendix: Discussion about baselines
Appendix C Appendix: More visualization results
@@
Appendix D Appendix: Future work
Moving forward, several possibilities for future research emerge. First, we aim to explore composition-aware fine-tuning for VLMs, which could potentially enable more effective utilization of multimodal information. Moreover, we are intrigued by the prospect of combining parameter-efficient tuning (He et al., 2022) with Helip  potentially further enhancing performance. We also intend to investigate how the integration of our boosting algorithm might alter the multimodal dataset curation algorithm (Gadre et al., 2023). Additionally, we propose extending our approach to the other contrastive learning domain, such as the graph contrastive learning (Wang et al., 2022a, b). Ultimately, we hope our work will serve as a catalyst for additional research in the fine-tuning of pre-trained, large-scale multimodal models.

### v3

@@ Appendix E Appendix: Performance of HELIP on noisy dataset
Appendix F Appendix: Analysis of the Impact of $\tau$ on Hard Pair Selection
To examine the impact of the threshold parameter $\tau$ on the selection of hard pairs, we analyze the similarities in the rankings of hard pairs (using Kendall Rank Similarity) mined by HPM under various $\tau$ values. The hard pairs are ranked by using the selection criteria value mentioned in Appendix C. The results on the CC12M dataset are displayed in Figure 7. We observe that the selection of hard pairs is robust to changes in the $\tau$ value. This resilience is partly because we only mine the top 50 hard pairs, a subset unlikely to be significantly affected when $\tau\leq 0.5$.
Figure 7: The Impact of $\tau$ on Hard Pair Selection.
Appendix G Appendix: Analysis of the Impact of Mitigating Noisy Data
As presented in Section 3.2, to enhance the overall quality and reliability of the training dataset, data pairs lacking substantial support from the entirety of the training data are considered unsuitable and removed.
This section further empirically analyzes the impact of our noise mitigation strategy by detailing the quantity and nature of pairs removed across various datasets. Specifically, our approach removes 4.67% of the pairs from CC3M, 3.64% from CC12M, and 7.41% from YFCC15M, before continuing with pretraining. Figure 8 visualizes the pairs filtered from CC12M. Notably, our strategy effectively removed pairs such as unavailable images (e.g., two blank or white images in the second row) and mismatched pairs. These results suggest that our noise mitigation strategy can effectively clean the data using two single-modality models before training a CLIP model from scratch.
Figure 8: Visualization of the image-caption pairs filtered out from CC12M.
@@ Appendix G Appendix: Analysis of the Impact of Mitigating Noisy Data
Appendix H Appendix: Comparison with Other Hard Data Selection Method
@@
We evaluate the efficacy of the proposed method in enhancing the discriminative capacity of learned representations by comparing its zero-shot classification performance with that of other hard data mining strategies. As described in the Section 2, a common way to define hard data is through intra-modality similarity. Hence, we introduce the hard data mining methods depending on (sample level) image similarity mining and text similarity mining and denote them as IM and TM respectively. For a given target pair, we compute the cosine similarity between its image/text representation and that of the remaining dataset. The image and text representations are encoded using a pretrained Resnet50 and BERT, respectively. As the preprocessing step, IM and TM methods mine hard negatives before continuous pretraining. Subsequently, we integrate the mined hard negative pairs into the training pipeline of CLIP and denote them as CLIP+IM and CLIP+TM and optimize the original contrastive loss to fine-tune the model. Additionally, we also include the hard negative contrastive loss, HN-NCE, proposed by  Radenovic et al. (2023), as a baseline. HN-NCE upsamples the weight of hard-negatives identified by the current model. As shown in Table 9, when the CC3M pretrained CLIP model is combined with Helip, the performance of our pair-level hard data mining method significantly outperforms other sample-level techniques. Besides,we observe that compared to the baseline CLIP performance, the introduction of TM and IM methods results in a decline in performance. To better understand the reasons behind this drop, we analyzed the outputs of the TM and IM methods in detail. In Figure9, we illustrate the data obtained through three distinct preprocessing methods: Hard Pair Mining (HPM), Image Similarity Mining (IM), and Text Similarity Mining (TM). The first row depicts the image-text pairs identified by HPM, while the second and third rows showcase the pairs mined by IM and TM, respectively. For TM (IM displays similar issues), the selected pairs often feature captions that are highly similar or identical, which is typical in data collected from the web. Even though identical pairs may not always be present, repetitions of the same images or text are common. According to the CLIP contrastive loss (Equation 1), the model is forced to push nearly identical caption representations toward and away from two distinct image representations at the same time.This inherent contradiction in objectives contributes to a degradation in performance. To illustrate, consider a target pair $(T_{\text{target}},I_{\text{target}})$ and a mined pair $(T_{\text{mined}},I_{\text{mined}})$ using TM, where $T_{\text{target}}\approx T_{\text{mined}}$ but $I_{\text{target}}\not\approx I_{\text{mined}}$. In the contrastive loss framework, the model aims to minimize the distance between $(I_{\text{target}},T_{\text{target}})$ and maximize the distance between $(I_{\text{target}},T_{\text{mined}})$. However, the near-identity of $T_{\text{target}}$ and $T_{\text{mined}}$ leads to conflicting optimization targets and a potential decline in performance.
@@
Appendix I Appendix: Impact of different encoders in HPM
We explored the effect of different pretrained encoders on HPM’s performance by alternating image and text encoders. Initially, the unsupervised pretrained DINO VITs8 (Caron et al., 2021) was paired with the SentenceT (Reimers and Gurevych, 2019) transformer, trained on over a billion internet-based sentences. This combination was later swapped for the SWAG VITb16 (Singh et al., 2022) and the T5 (Raffel et al., 2020). Additionally, experiments using OpenAI’s CLIP model (Radford et al., 2021) multimodal encoders were conducted. Interestingly, as Table 10 suggests, the encoder choice seemingly has negligible impact on HPM’s performance, likely due to the proficiency of current pretrained models in modeling intra-modal similarities. Moreover, the ability to use single-modal pretrained models and still achieve competitive or superior performance implies that there’s no assumption of having access to a high-quality CLIP model, such as OpenAI’s CLIP-400M.
@@ Appendix I Appendix: Impact of different encoders in HPM
Appendix J Appendix: More visualization results
@@
Appendix K Appendix: Future work
Moving forward, several possibilities for future research emerge. First, we aim to explore composition-aware fine-tuning for VLMs, which could potentially enable more effective utilization of multimodal information. Moreover, we are intrigued by the prospect of combining parameter-efficient tuning (He et al., 2022) with Helip  potentially further enhancing performance. Another area of interest is scaling up the dataset size and examining the applicability of the scaling law to our method. We also intend to investigate how the integration of our boosting algorithm might alter the multimodal dataset curation algorithm (Gadre et al., 2023). Ultimately, we hope our work will serve as a catalyst for additional research in the fine-tuning of pre-trained, large-scale multimodal models.
Appendix L Appendix: Potential Risks and Ethical Considerations
While our research contributes to advancements in language-image pretraining, it is important to acknowledge potential risks and ethical considerations associated with our work.
Privacy Concerns: The datasets used—CC3M, CC12M, subsets of YFCC15M, LAION7.5M, and LAION8M—are publicly available and sourced from the internet. They may contain personal identifiable information (PII) or images of individuals. We rely on the dataset providers’ curation processes to remove personal identifiable information.
Bias and Fairness Issues: The datasets may not be representative of all demographics or cultures, potentially leading to models that perform unevenly across different groups. This lack of fairness can perpetuate existing societal biases and inequalities.
Misuse of Technology: The models developed could be misused for malicious purposes, such as generating deepfakes, enabling unauthorized surveillance, or creating misleading information, which could have negative societal impacts.
To address these potential risks, the following steps could be considered in future work:
Enhanced Data Filtering: Currently, we rely on the curation and filtering processes conducted by the dataset creators for the datasets we used—CC3M, CC12M, YFCC15M subsets, and subsets of LAION. Implementing additional data cleaning procedures to identify and remove personally identifiable information (PII) and offensive content from the training datasets is an important area deserving future study.
Bias Mitigation Techniques: Incorporate fairness-aware learning algorithms and conduct thorough evaluations to detect and reduce biases in the model’s output.
Transparency and Accountability: Provide documentation detailing the data sources, model limitations, and potential biases to inform users and stakeholders.
Appendix M Appendix: Licensing
All datasets used in this study are open-source and utilized in accordance with their respective licenses. Specifically, CC3M and CC12M comprise images with Creative Commons licenses, YFCC100M is distributed under a Creative Commons Attribution license, and LAION-5B consists of web-sourced data under permissive licenses. Our use of these datasets is strictly for research purposes and complies with their licensing terms.
Appendix N Appendix: Use of AI Assistants
We acknowledge the use of AI assistants in the preparation of this work. Specifically, we utilized tools, GPT4-o, for proofreading the manuscript and GitHub Copilot for assisting with coding tasks. These AI tools were employed to enhance productivity and efficiency. All content generated with the assistance of AI was thoroughly reviewed and edited by the authors to ensure accuracy and originality. The responsibility for the final content of this paper rests solely with the authors.
Appendix O Appendix: Artifact Use Consistent with Intended Purpose
In our research, we utilized several publicly available datasets—CC3M, CC12M, YFCC15M subsets, LAION7.5M, and LAION8M—that are widely used within the computer vision and machine learning communities for the purpose of training and evaluating language-image models. The use of these datasets in our study is consistent with their intended purpose, as specified by their creators, which is to advance research in image recognition, captioning, and related fields. All datasets were used strictly for non-commercial, research-oriented objectives, adhering to the access conditions and licenses provided by the dataset providers.
For the artifacts we created during this research, including the trained models and any derived datasets or code, we specify that their intended use is for academic and research purposes only. These artifacts are shared to promote transparency, reproducibility, and further advancement in the field. We ensure that this intended use is compatible with the original access conditions of the datasets we used, particularly considering any restrictions on derivative works or redistribution. When sharing our artifacts, we comply with all applicable licenses and access terms, and we encourage others who use our artifacts to do the same.